OpenStudy (anonymous):

HHHHHHHHEEEEEEEEELLLLLPPPPPPPPPP 8^(-x-2)=237

OpenStudy (anonymous):

\[8^{(-x-2)}=237\]

OpenStudy (anonymous):

Take the log8 of both sides.\[\log_{8} 8^{(-x-2)}=\log_{8} 237\] \[-x-2=\log_{8} 237\]\[-x=\log_{8} +2\]\[x=\frac{ \log_{8}+2 }{ -1 }\] Change the base of log8 \[x=\frac{ -\ln 237 }{ \ln 8 }-2\]